# Efficient and high-performance routing of lattice-surgery paths on three-dimensional lattice

### v3

@@ 4.4 Discussion
First, we discuss characteristics seen both in random circuits and the SELECT circuits. For throughput, BFS, look-ahead BFS, 3D BFS, and 3D Dijkstra performed well in that order, with Dijkstra projection performing as well as 3D Dijkstra. Look-ahead optimization in Dijkstra projection had no impact on random circuits while it proved effective in SELECT circuits. It is natural that look-ahead BFS outperforms BFS, considering that look-ahead BFS executes future operations and produces more spare space that would be blocked in BFS. The primary reason 3D BFS outperforms look-ahead BFS is its ability to pack paths more densely into the lattice by increasing the dimension of the paths by one. Another reason is that a simple look-ahead can be reproduced by searching in 3D. The reason 3D Dijkstra outperforms 3D BFS is that the paths can be packed more densely into the lattice in the end by keeping the overall path low, even if it is a bit roundabout, rather than unnecessarily shortening the path by going through an excessively high layer. The reason Dijkstra projection performs comparably to 3D Dijkstra is likely that the paths adopted by 3D Dijkstra tend to be along the bottom surface, as in Dijkstra projection.
In terms of execution time, BFS was the shortest, look-ahead BFS was slightly longer than BFS, Dijkstra projection was a few times longer, and 3D BFS and 3D Dijkstra were much longer than the others. From these results, execution time is significantly affected by whether the search space is a 2D or 3D lattice. The fact that look-ahead BFS is slightly slower than BFS is a natural consequence of the additional cost of instruction look-ahead. The reason Dijkstra projection is several times slower than BFS and look-ahead BFS is that, as a comparison of search algorithms, Dijkstra’s algorithm is more computationally expensive than BFS. The same argument holds for 3D BFS and 3D Dijkstra.
@@ 4.5 Resource overhead analysis
The quantitative compilation-time overhead was shown in the previous section as execution time. The results indicate that employing the Dijkstra projection with kink constraints increases the compilation time per instruction by a factor of a few. For the SELECT instances, which are large enough to demonstrate quantum advantage [yoshioka2022hunting], the compilation completes in approximately half a minute. This duration is negligible compared to the total execution time, as the quantum phase estimation with the corresponding SELECT circuit on a real FTQC would take over an hour [yoshioka2022hunting]. Even for larger instances, both the compilation time with Dijkstra projection and the execution time of an actual machine scale almost linearly with the number of instructions, as summarized in Tab. 2. Therefore, the use of the Dijkstra projection introduces no practical overhead in regimes where quantum computation is meaningful.
@@
A potential overhead in using the Dijkstra projection is an increase in the error rate. Let the number of voxels in the 3D representation be denoted as the active volume. Because 3D routing might entail longer routing paths than 2D routing, the active volume can be larger. We denote the probability of maintaining a single logical cell fault-tolerantly as $p_{\rm voxel}$ and the active volume as $A$. Then, the overall failure probability can be approximated as $p_{\rm voxel}A$. Hence, a larger active volume increases the probability of program failure. We evaluated the increase in active volume using 3D-routing-based methods for the SELECT circuits, as shown in Tab. 7. The results show that, for the SELECT 0 instance, the look-ahead Dijkstra projection with kink condition increases the active volume by about 30% compared to BFS without kink conditions. Conversely, the volume decreases for the SELECT 5 instance. We expect this is because improved instruction throughput due to 3D routing reduces the idle time of logical qubits. As discussed below, the total program failure rate can be kept constant by increasing the qubit count, and a 30% increase in active volume can be compensated with a modest cost.
@@
Next, we examine the qubit-count penalty. The total number of physical qubits in an FTQC system is the product of the number of logical cells in the qubit plane and the number of physical qubits per logical cell. In surface codes, the latter is $2d^{2}-1$, where $d$ is the code distance. The introduction of 3D routing does not increase the number of logical cells. Although 3D routing temporarily allocates logical qubits to separate logical instructions, they are assigned to existing logical cells. In contrast, the number of physical qubits per logical cell may increase. If the active volume increases, the error-correction failure rate per voxel $p_{\rm voxel}$ must be reduced to maintain a constant overall program failure probability, which requires increasing the code distance. The error-estimation failure probability per syndrome-measurement cycle on a single logical cell is approximated as $0.1(100p)^{(d+1)/2}$ [litinski2019game], where $p$ is the physical error rate per cycle. Consequently, the failure probability per voxel (per $d$ cycles) is $p_{\rm voxel}=0.1d(100p)^{(d+1)/2}$. Assuming $p=10^{-3}$ and $d$ is sufficiently large, increasing (decreasing) the code distance by two allows for roughly 10 times larger (smaller) active volume. In practical regimes, the code distance $d$ is typically 23 or higher [yoshioka2022hunting]. Increasing $d$ from 23 to 25 corresponds to approximately an 18% increase in total physical qubits. As discussed earlier, 3D routing can either increase or decrease the active volume by a few times, which corresponds to at most an 18% variation in total physical qubits.
@@
Table 7: Active volume evaluation on SELECT circuits. The grayed rows annotated with an asterisk represent reference values that disregard the kink condition and assume CNOT operations complete via a 2D path in $1$ code beat. Although this assumption makes their outputs infeasible, doubling these values will provide the corresponding feasible values. The remaining methods adhere to the kink condition and yield a feasible solution.
@@ 5 Conclusion and future outlook
In this paper, we proposed efficient and high-performance algorithms for lattice-surgery instructions. We showed that the mapping of lattice-surgery instructions can be reduced to 3D path search problems. Based on this mapping, we proposed Dijkstra projection and its look-ahead variant to perform efficient scheduling of lattice-surgery instructions. We evaluated their performance on several settings and they showed significant improvements in throughput compared to the baseline while the increase in execution time is modest.
@@
In our numerical experiment, only the Pauli-$XX$ and $ZZ$ measurements were considered, and other logical operations were ignored for simplicity. However, we need to consider other types of instructions in practice, such as single-qubit Clifford gates, lattice-surgery instructions acting on many logical-qubit cells, and magic-state teleportations. In particular, many-body lattice-surgery operations are vital since they are used to remove Clifford unitary operations from the circuits [litinski2019game] and used in several existing works [litinski2019game, silva2024multi]. In Appendix A, we provide the extension of our techniques to many-body lattice surgery.
@@
One may also be concerned that logical qubit cells are blocked during the reaction time in magic-state teleportation. Typical $T$-gate teleportation circuits require $S$-gate feedback according to the measurement results, and we cannot perform operations on cells that may be targets of the feedback during the latency. Nevertheless, we expect this is not a problem since we can relax this blocking by using auto-corrected $T$-gate teleportation [litinski2019game, fowler2012time], which converts feedback operations from $S$-gate on the computing space to the choice of Pauli measurement basis on the ancillary space. Combining these theoretical ideas and providing consistent and full-stack compilation frameworks such as Refs. [beverland2022assessing, zhu2024ecmas, zhu2025ecmas+, watkins2023high], would be the next direction of this topic. To illustrate the contexts intended by our method, we describe an example full-stack compilation workflow in Appendix B.
@@
There are several other directions for future work. For the voxel weighting in 3D Dijkstra and Dijkstra projection, we proposed the weighting that worked best among those we tried, but we expect that there is room for further improvement. Also, the theoretical upper and lower bounds of the idea, separating large instructions into small fragments, may be calculated in more detail in future. By comparing these bounds with our numerical results, we can measure the room for improvement. While we introduced a method to satisfy the restriction of kink parity for lattice-surgery instructions, there may be a more efficient way to control the number of kinks. Integrating kink restrictions in the path-finding algorithms more naturally would be another future topic.
@@
Integrating our method into distributed FTQC would also be essential. The integration method strongly depends on the paradigm of distributed FTQC. If the number of qubits per node is sufficiently small, logical qubits are constructed through communication among several nodes [nickerson2013topological] or realized via measurement-based FTQC [bombin2021interleaving, bourassa2021blueprint]. In such cases, the routing problem is equivalent to that in the single-node scenario, and our method can be applied without modification. The situation becomes more complex when the number of qubits per node is moderate. In this regime, each node can host several logical qubits, while two-qubit logical operations are performed via logical entanglement distribution and teleportation [pattison2025constant, maeda2025logical]. Since this inter-node teleportation can be regarded as a sparse connection between 3D voxels, we believe our method can be extended to such cases. Nevertheless, the detailed implementation of distributed FTQC architectures remains immature and will be an important subject for future work.
@@ Appendix A Extension to many-body lattice surgery
The proof and algorithm presented in the main text focused on 3D routing for two-body lattice surgery. Two-body lattice surgery is a crucial case because quantum circuits are typically expressed in the Clifford+T form, where S and H gates are executed via code deformation [brown2017poking, fowler2018low], and CNOT and T gates are translated into two-body lattice surgeries [chamberland2022universal]. On the other hand, all Clifford gates can be omitted by converting two-body lattice surgeries into many-body ones, provided there is no intermediate measurement [litinski2019game]. It is therefore important to apply the key idea of this paper, decomposing a large lattice surgery into smaller fragments, to such cases. In this section, we show that the concept of this paper can be straightforwardly extended to many-body cases.
@@ A.1 Conditions for 3D routing for many-body lattice surgery
Condition 2 (No-temporal-fork): No voxel is connected to the top or bottom of a fork voxel. Thus, Fig. 16 is a fork satisfying this condition while Figs. 16 and 16 do not.
@@
(a) Example of segment separation
(b) Corresponding circuits
@@
Suppose there are $m$ logical qubit cells. A 3D routing satisfying the following conditions can be converted into $m$-body Pauli-ZZ($X$) measurements on $m$ logical qubits.
@@
Note that if we want to perform lattice surgery in which the Pauli operator contains both $X$ and $Z$, we can perform it by adjusting the position of the surface-code lattice in advance. This requires the ability to shift or rotate the shape of surface codes and to perform stabilizer Pauli measurements that involve both $X$ and $Z$ [fowler2018low, litinski2019game, mcewen2023relaxing], which would incur additional latency in scheduling.
@@ A.2 Proof of theorem
  • Every odd segment forms a path between $X$-boundaries of two ancillary cells.
@@
Next, we reduce a sequence consisting of an even segment, an ancillary cell, an odd segment, another ancillary cell, and an even segment into a single even segment using the circuit equivalence shown in Figs. 18, 18, 18. This procedure is repeated until no ancillary cells or odd segments remain. Eventually, a single even segment connects all the leaves of the original tree, which corresponds to the target many-body Pauli measurement.
@@
Figure 18: Equivalent processes of many-body measurements. Here, $P$ and $Q$ are an arbitrary many-body Pauli operator, and $M_{AB}$ is a Pauli measurement with projectors $(I\pm A\otimes B)/2$ on two registers. (a) Two many-body measurements that act on ancillary qubits with the same basis are equal to a single many-body measurement. (b,c,d) Three alternating many-body measurements are equal to a single many-body measurement. These are many-body variants of Fig. 7.

### v4

@@ 4.4 Discussion
First, we discuss characteristics seen both in the random circuits and the SELECT circuits. For throughput, the BFS, the look-ahead BFS, the 3D BFS, and the 3D Dijkstra performed well in that order, with the Dijkstra projection performing as well as the 3D Dijkstra. The look-ahead optimization in the Dijkstra projection had no impact on the random circuits, while it proved effective in SELECT circuits. It is natural that the look-ahead BFS outperforms the BFS, considering that the look-ahead BFS executes future operations and produces more spare space that would be blocked in the BFS. The primary reason the 3D BFS outperforms the look-ahead BFS is its ability to pack paths more densely into the lattice by increasing the dimension of the paths by one. Another reason is that a simple look-ahead can be reproduced by searching in 3D. The reason the 3D Dijkstra outperforms the 3D BFS is that the paths can be packed more densely into the lattice in the end by keeping the overall path low, even if it is a bit roundabout, rather than unnecessarily shortening the path by going through an excessively high layer. The reason the Dijkstra projection performs comparably to the 3D Dijkstra is likely that the paths adopted by the 3D Dijkstra tend to be along the bottom surface, as in the Dijkstra projection.
In terms of execution time, the BFS was the shortest, the look-ahead BFS was slightly longer than the BFS, the Dijkstra projection was a few times longer, and the 3D BFS and the 3D Dijkstra were much longer than the others. From these results, execution time is significantly affected by whether the search space is a 2D or 3D lattice. The fact that the look-ahead BFS is slightly slower than the BFS is a natural consequence of the additional cost of instruction look-ahead. The reason the Dijkstra projection is several times slower than the BFS and the look-ahead BFS is that, as a comparison of search algorithms, Dijkstra’s algorithm is more computationally expensive than the BFS algorithm. The same argument holds for the 3D BFS and the 3D Dijkstra.
@@ 4.5 Resource overhead analysis
The quantitative compilation-time overhead was shown in the previous section as execution time. The results indicate that employing the Dijkstra projection with kink constraints increases the compilation time per instruction by a factor of a few. For the SELECT instances, which are large enough to demonstrate quantum advantage [44], the compilation completes in approximately half a minute. This duration is negligible compared to the total execution time, as the quantum phase estimation with the corresponding SELECT circuit on a real FTQC would take over an hour [44]. Even for larger instances, both the compilation time with the Dijkstra projection and the execution time of an actual machine scale almost linearly with the number of instructions, as summarized in Tab. 2. Therefore, the use of the Dijkstra projection introduces no practical overhead in regimes where quantum computation is meaningful.
@@
A potential overhead in using the Dijkstra projection is an increase in the error rate. Let the number of voxels in the 3D representation be denoted as the active volume. Because 3D routing might entail longer routing paths than 2D routing, the active volume can be larger. We denote the probability of maintaining a single logical qubit fault-tolerantly for a single code beat as $p_{\rm voxel}$ and the active volume as $A$. Then, under the assumption that $p_{\rm voxel}$ is sufficiently small and that errors are largely independent, the overall failure probability can be approximated as $p_{\rm voxel}A$. Hence, the active volume serves as a semi-quantitative proxy of the overall failure probability, provided that the errors are approximately independent and that $p_{\rm voxel}$ is low enough to keep the overall failure risk within a practical range. We evaluated the increase in active volume using 3D-routing-based methods for the SELECT circuits, as shown in Tab. 7. As a conservative analysis, our proposed method satisfying the kink condition is compared with the baseline method disregarding the kink condition. The results show that, for the SELECT 0 instance, the look-ahead Dijkstra projection with kink condition increases the active volume by about 30% compared to the BFS without kink conditions. Conversely, the volume decreases for the SELECT 5 instance. We expect this is because improved instruction throughput due to 3D routing reduces the idle time of logical qubits. As discussed below, the total program failure rate can be kept constant by increasing the qubit count, and a 30% increase in active volume can be compensated with a modest cost.
@@
Next, we examine the qubit-count penalty. The total number of physical qubits in an FTQC system is the product of the number of logical-qubit cells in the qubit plane and the number of physical qubits per logical-qubit cell. In surface codes, the latter is $2d^{2}-1$, where $d$ is the code distance. The introduction of 3D routing does not increase the number of logical-qubit cells. Although 3D routing temporarily allocates logical qubits to separate logical instructions, they are assigned to existing cells. In contrast, the number of physical qubits per logical-qubit cell may increase. If the active volume increases, the error-correction failure rate per voxel $p_{\rm voxel}$ must be reduced to maintain a constant overall program failure probability, which requires increasing the code distance. The error-estimation failure probability per syndrome-measurement cycle on a single cell is approximated as $0.1(100p)^{(d+1)/2}$ [26], where $p$ is the physical error rate per cycle. Consequently, the failure probability per voxel (per $d$ cycles) is $p_{\rm voxel}=0.1d(100p)^{(d+1)/2}$. Assuming $p=10^{-3}$ and $d$ is sufficiently large, increasing (decreasing) the code distance by two allows for roughly 10 times larger (smaller) active volume. In practical regimes, the code distance $d$ is typically 23 or higher [44]. Increasing $d$ from 23 to 25 corresponds to approximately an 18% increase in total physical qubits. As discussed earlier, 3D routing can either increase or decrease the active volume by a few times, which corresponds to at most an 18% variation in total physical qubits.
@@
Table 7: Active volume evaluation on SELECT circuits. The grayed rows annotated with an asterisk represent reference values that disregard the kink condition, optimistically assuming that CNOT operations complete via a 2D path in one code beat. Although these raw infeasible values are presented to show their net routing performance, doubling these values will provide the corresponding feasible values, reflecting that CNOT operations realistically require two code beats. The remaining methods adhere to the kink condition and yield a feasible solution.
@@
4.6 Robustness under Resource Pressure
To evaluate the robustness of our proposed methods under high resource contention, we examined their performance across circuits with varying gate densities. We define the average gate density as the number of instructions divided by the circuit depth, which quantifies the congestion level of the problem instance. This metric also serves as a theoretical upper bound for the throughput of the 2D routing methods, BFS and look-ahead BFS. As discussed in Sec. 4.2, we utilized the SELECT circuits to generate different levels of congestion by varying the parallelism parameter from 0 to $8$: SELECT 0 corresponds to a serial implementation, and SELECT $i$ is parallelized into $2^{i}$ threads.
Figure 16: Degradation curves of throughput as a function of gate density. The performance is measured by the throughput normalized by the average gate density, which corresponds to an ideal throughput of conflict-free 2D routing.
Figure 17: Degradation curves of active volume as a function of gate density.
Fig. 16 illustrates the performance degradation of the scheduling methods as resource pressure intensifies. The performance is quantified by the throughput normalized by the average gate density, which is equivalent to the ratio of the circuit depth to the total execution time in code beats. To assess the impact of routing conflicts in the proposed methods, we also included the normalized performance of the conflict-free 3D routing, that reserves only path endpoints and ignores all intermediate segments. This benchmark serves as a theoretical upper bound on the performance of 3D routing methods.
As expected, all methods exhibit a decline in efficiency as gate density increases. Notably, at low gate densities, the proposed methods achieve near-optimal performance, outperforming the circuit depth. This can be attributed to their advantages in stair-like structures, as discussed in Sec. 3.4. Increasing gate density leads to performance degradation across all routing methods, resulting in a growing discrepancy compared to the theoretical 3D upper bound due to intensified path conflicts. However, it is noteworthy that the proposed methods remain comparable to the circuit depth, demonstrating their resilience to high resource pressure.
Fig. 17 shows the active volume of the routing methods in the same numerical experiment. While all methods experience degradation, the proposed methods maintain an active volume close to that of the look-ahead BFS across different gate densities. Furthermore, they outperform the look-ahead BFS in higher densities due to their shorter execution time in code beat.
@@ 5 Conclusion and future outlook
In this paper, we proposed efficient and high-performance algorithms for lattice-surgery instructions. We showed that the mapping of lattice-surgery instructions can be reduced to 3D path search problems. Based on this mapping, we proposed the Dijkstra projection and its look-ahead variant to perform efficient scheduling of lattice-surgery instructions. We evaluated their performance on several settings, and they showed significant improvements in throughput compared to the baseline, while the increase in execution time is modest.
@@
In our numerical experiment, only the Pauli-$XX$ and $ZZ$ measurements were considered, and other logical operations were ignored for simplicity. However, we need to consider other types of instructions in practice, such as single-qubit Clifford gates, lattice-surgery instructions acting on many logical-qubit cells, and magic-state teleportations. In particular, many-body lattice-surgery operations are vital since they are used to remove Clifford unitary operations from the circuits [26] and used in several existing works [26, 39]. In Appendix A, we provide the extension of our techniques to many-body lattice surgery.
@@
One may also be concerned that logical-qubit cells are blocked during the reaction time in magic-state teleportation. Typical $T$-gate teleportation circuits require $S$-gate feedback according to the measurement results, and we cannot perform operations on cells that may be targets of the feedback during the latency. Nevertheless, we expect this is not a problem since we can relax this blocking by using auto-corrected $T$-gate teleportation [26, 15], which converts feedback operations from $S$-gate on the computing space to the choice of Pauli measurement basis on the ancillary space.
Combining these theoretical ideas and providing consistent and full-stack compilation frameworks, such as Refs. [3, 47, 46, 41], would be the next direction of this topic. To illustrate the contexts intended by our method, we describe an example full-stack compilation workflow in Appendix B. As explained in this appendix, our method is compatible with lattice-surgery-based FTQC pipelines under the following four assumptions. (1) The positions and connectivity of physical qubits are fixed and cannot be changed during computation. (2) Logical qubits are encoded using surface codes, and logical operations are implemented via code deformation and lattice surgery. (3) All required surface-code logical qubits are allocated at the beginning of execution and are not deallocated during program execution. (4) Magic states are generated and managed within abstracted modules placed at fixed locations, and we ignore probabilistic behavior at runtime for simplicity. We believe that these assumptions are common when evaluating architectures with superconducting qubits [3, 19, 18], and the third and fourth assumptions would be removed as they are assumed to simplify the compilation and evaluation process. On the other hand, the other two assumptions do not necessarily hold when integrating recent technologies such as qubit shuttling in neutral-atom platforms [45], non-topological or concatenated quantum error-correcting codes [17, 43, 42], and generalized lattice-surgery operations for such codes [10]. Extending our method to a wider range of compilation frameworks and evaluating the performance of our methods in broader situations are expected, which are left for future work.
@@
There are several other directions for future work. For the voxel weighting in the 3D Dijkstra and the Dijkstra projection, we proposed the weighting that worked best among those we tried, but we expect that there is room for further improvement. Also, the theoretical upper and lower bounds of the idea, separating large instructions into small fragments, may be calculated in more detail in the future. By comparing these bounds with our numerical results, we can measure the room for improvement. While we introduced a method to satisfy the restriction of kink parity for lattice-surgery instructions, there may be a more efficient way to control the number of kinks. Integrating kink restrictions in the path-finding algorithms more naturally would be another future topic.
@@
Integrating our method into distributed FTQC would also be essential. The integration method strongly depends on the paradigm of distributed FTQC. If the number of qubits per node is sufficiently small, logical qubits are constructed through communication among several nodes [33] or realized via measurement-based FTQC [5, 6]. In such cases, the routing problem is equivalent to that in the single-node scenario, and our method can be applied without modification. The situation becomes more complex when the number of qubits per node is moderate. In this regime, each node can host several logical qubits, while two-qubit logical operations are performed via logical entanglement distribution and teleportation [35, 30]. Since this inter-node teleportation can be regarded as a sparse connection between 3D voxels, we believe our method can be extended to such cases. Nevertheless, the detailed implementation of distributed FTQC architectures remains immature and will be an important subject for future work.
@@ Appendix A Extension to many-body lattice surgery
The proof and algorithm presented in the main text focused on 3D routing for two-body lattice surgery. Two-body lattice surgery is a crucial case because quantum circuits are typically expressed in the Clifford+T form, where S and H gates are executed via code deformation [8, 12], and CNOT and T gates are translated into two-body lattice surgeries [9]. On the other hand, all Clifford gates can be omitted by converting two-body lattice surgeries into many-body ones, provided there is no intermediate measurement [26]. It is therefore important to apply the key idea of this paper, decomposing a large lattice surgery into smaller fragments, to such cases. In this section, we show that the concept of this paper can be straightforwardly extended to many-body cases.
@@ A.1 Conditions for 3D routing for many-body lattice surgery
Condition 2 (No-temporal-fork): No voxel is connected to the top or bottom of a fork voxel. Thus, Fig. 18 is a fork satisfying this condition while Figs. 18 and 18 do not.
@@
(a) Example of segment separation.
(b) Corresponding circuits.
@@
Suppose there are $m$ target data cells. A 3D routing satisfying the following conditions can be converted into $m$-body Pauli-ZZ($X$) measurements on $m$ target cells.
@@
Note that if we want to perform lattice surgery in which the Pauli operator contains both $X$ and $Z$, we can perform it by adjusting the position of the surface-code lattice in advance. This requires the ability to shift or rotate the shape of surface codes and to perform stabilizer Pauli measurements that involve both $X$ and $Z$ [12, 26, 31], which would incur additional latency in scheduling.
@@ A.2 Proof of theorem
  • Every odd segment forms a path between the $X$-boundaries of two ancillary cells.
@@
Next, we reduce a sequence consisting of an even segment, an ancillary cell, an odd segment, another ancillary cell, and an even segment into a single even segment using the circuit equivalence shown in Figs. 20, 20, 20. This procedure is repeated until no ancillary cells or odd segments remain. Eventually, a single even segment connects all the leaves of the original tree, which corresponds to the target many-body Pauli measurement.
@@
Figure 20: Equivalent processes of many-body measurements. Here, $P$ and $Q$ are an arbitrary many-body Pauli operator, and $M_{AB}$ is a Pauli measurement with projectors $(I\pm A\otimes B)/2$ on two registers. (a) Two many-body measurements that act on an ancillary logical qubit with the same basis are equal to a single many-body measurement. (b,c,d) Three alternating many-body measurements are equal to a single many-body measurement. These are many-body variants of Fig. 7.